# Isoparametric foliations and bounded geometry

### v1

@@ 4. Examples of infinite families of foliations
induced by restricting to the diffeomorphisms on the base. For $m\geq 1$, we have $\pi_{0}(\operatorname{Map}(\mathbb{S}^{m},\mathrm{O}(k)))=\pi_{m}(\mathrm{O}(k))\rtimes\mathbb{Z}/2$, so as $\pi_{m}(O(2))=\pi_{m}(\mathbb{S}^{1})=0$ for $m\geq 2$ and $\pi_{1}(O(m))=\mathbb{Z}/2$ for $m\geq 3$, we see that the two mapping spaces in (4.1) have finitely many components. Since $\pi_{0}(\operatorname{Diff}(\mathbb{S}^{m}))$ is finite for $m\neq 4$ by Corollary 5.2 below, it follows from the long exact sequence in homotopy groups that $\pi_{0}(\operatorname{Isom}_{b}(D^{\pm}))$ and thus $H_{\pm}$ is finite for $n\neq 6$. For $n=6$, the group $\pi_{0}(\operatorname{Diff}(\mathbb{S}^{n-2}))$ maybe be infinite, but since $\pi_{0}(\operatorname{Map}(\mathbb{S}^{n-2},O(2)))$ is finite even for $n=6$ and $\partial D^{\pm}=\mathbb{S}^{n-2}\times\mathbb{S}^{1}$, for finiteness of $H_{+}$ it actually suffices that that the image of ()×$(-)\times\mathbb{S}^{1}\colon\pi_{0}(\operatorname{Diff}(\mathbb{S}^{n-2}))\rightarrow\pi_{0}(\operatorname{Diff}(\mathbb{S}^{n-2}\times\mathbb{S}^{1}))$ is finite for $n=6$ which holds by Lemma 5.3 below. ∎
@@
The space of leaves $X_{m}$ of $\mathcal{F}_{m}$ is canonically isometric to the quotient of $\tilde{M}$ by the action of the group (×$\mathbb{(}\mathbb{Z}\times H_{1})\cdot\mathcal{H}_{m}$. Hence, $X_{m}$ is the circle of length $\frac{1}{m}$. Thus, the leaves of $\mathcal{F}_{m}$ are fibers of a Riemannian submersion $P_{m}:M\to X_{m}$. As explained above, its fibers have fundamental groups $H_{m}$ and constant mean curvature, and this mean curvature does not depend on the fiber. Taking the fiber with minimal volume, we deduce that at this fiber the mean curvature is 00, thus all fibers of $P_{m}$ are minimal. Composing $P_{m}$ with an isoparametric function $g_{m}:X_{m}\to\mathbb{R}$, we obtain an isoparametric function $f_{m}:=g_{m}\circ P_{m}:M\to\mathbb{R}$ by [GT13, Proposition 2.4]. The isoparametric foliation induced by $f_{m}$ is exactly $\mathcal{F}_{m}$. But the fundamental groups of the fibers of $\mathcal{F}_{m}$ are different, for different $m$. This implies the claim for $\dim(M)=4$ and thus also in higher dimensions by taking products with a torus. ∎
@@ 5. Topological background
In this final section, we collect the differential-topological results used in the previous sections. To high-dimensional topologists, essentially all of them are either clear or well-known; we merely explain how to extract them from the literature for more geometrically minded readers. We adopt the notation from Section 2.3. In particular, for closed manifolds $M$ we write $\operatorname{Diff}(M)=\operatorname{Diff}(M,M)$ and $\operatorname{Homeo}(M)=\operatorname{Homeo}(M,M)$ for the topological groups of diffeomorphisms or homeomorphisms of $M$. If $M$ is compact but may have boundary, we require diffeomorphisms or homeomorphisms to agree with the identity in a neighbourhood of the boundary, which we indicate by a $\partial$-subscript. If we restrict to orientation-preserving diffeomorphisms or homeomorphisms, we add a $+$-superscript.
@@
By the parametrised isotopy extension sequence (see e.g. [Wal16, Theorem 6.1.1]), restriction along $D^{n}_{-}\subset\mathbb{S}^{n}$ yields a fibration $\operatorname{Diff}^{+}(\mathbb{S}^{n})\rightarrow\mathrm{Emb}^{+}(D^{n}_{-},\mathbb{S}^{n})$ to the space of orientation-preserving embeddings $D^{n}_{-}\hookrightarrow\mathbb{S}^{n}$. The fiber over the inclusion is $\operatorname{Diff}_{\partial}(D_{+}^{n})$. Taking derivatives induces a homotopy equivalence $\mathrm{Emb}^{+}(D^{n}_{-},\mathbb{S}^{n})\simeq\mathrm{GL}^{+}_{n+1}(\mathbb{R})\simeq\mathrm{SO}(n+1)$, so the standard $\mathrm{SO}(n+1)$-action on $\mathbb{S}^{n}$ yields a section of $\operatorname{Diff}^{+}(\mathbb{S}^{n})\rightarrow\mathrm{Emb}^{+}(D^{n}_{-},\mathbb{S}^{n})$ up to homotopy. The long exact sequence in homotopy groups then implies that $\mathrm{ext}_{n}$ is an isomorphism, so $\smash{\widetilde{\mathrm{ext}}_{n}}$ is surjective too. To show injectivity of $\widetilde{\mathrm{ext}}_{n}$, we prove that $\widetilde{\mathrm{ext}}_{n}\circ\mathrm{glue}_{n}$ is injective: let $[\varphi]$ be a class in the kernel, then there is an orientation-preserving diffeomorphism
$\smash{D^{n+1}_{+}\cup_{\mathrm{glue}(\varphi)}(\mathbb{S}^{n}\times[0,1])\cup_{\mathbb{S}^{n}\times\{1\}}D^{n+1}_{-}\overset{\Phi}{\cong}D^{n+1}_{+}\cup_{\mathrm{id}}(\mathbb{S}^{n}\times[0,1])\cup_{\mathbb{S}^{n}\times\{1\}}D^{n+1}_{-}.}$
@@
Since $\mathrm{ext}_{4}$ from (5.1) is an isomorphism by Lemma 5.1, it suffices to show that the composition $\pi_{0}(\operatorname{Diff}_{\partial}(D_{+}^{4}))\rightarrow\pi_{0}(\operatorname{Diff}_{\partial}(D_{+}^{4}\times\mathbb{S}^{1}))\rightarrow\pi_{0}(\operatorname{Diff}^{+}(\mathbb{S}^{4}\times\mathbb{S}^{1}))$ is trivial. We will see that this is already the case for the first map. Fix a basepoint $*\in D^{1}_{+}\subset\mathbb{S}^{1}$ and consider the restriction map $\pi_{0}(\operatorname{Diff}_{\partial}(D_{+}^{4}\times\mathbb{S}^{1}))\rightarrow\pi_{0}(\mathrm{Emb}_{\partial}(D_{+}^{4}\times*,D_{+}^{4}\times\mathbb{S}^{1}))$ to the set of isotopy classes of embeddings that agree with the inclusion $\mathrm{inc}\colon D_{+}^{4}\times*\subset D_{+}^{4}\times\mathbb{S}^{1}$ near the boundary. By isotopy extension, the fibers of this map are either empty or admit a surjection from $\pi_{0}(\operatorname{Diff}_{\partial}(D_{+}^{4}\times D^{1}_{-}))\cong\pi_{0}(\operatorname{Diff}_{\partial}(D^{5}))$ which is trivial by Lemma 5.1, so $\pi_{0}(\operatorname{Diff}_{\partial}(D_{+}^{4}\times\mathbb{S}^{1}))\rightarrow\pi_{0}(\mathrm{Emb}_{\partial}(D_{+}^{4}\times*,D_{+}^{4}\times\mathbb{S}^{1}))$ is injective. It thus suffices to show that the map $\pi_{0}(\operatorname{Diff}_{\partial}(D_{+}^{4}))\rightarrow\pi_{0}(\mathrm{Emb}_{\partial}(D_{+}^{4}\times*,D_{+}^{4}\times\mathbb{S}^{1}))$ which sends $[\varphi]\in\pi_{0}(\operatorname{Diff}_{\partial}(D_{+}^{4}))$ to $[\mathrm{inc}\circ\varphi]$ is constant. By identifying $D_{+}^{1}\subset\mathbb{S}^{1}$ with $[0,1]$ one sees that this maps sends concordant diffeomorphisms to the same class of embeddings, so it factors over the group $\widetilde{\pi}_{0}(\operatorname{Diff}_{\partial}(D_{+}^{4}))$. The latter is trivial by Lemma 5.1, so the claim follows. ∎

### v2

@@ 4. Examples of infinite families of foliations
induced by restricting to the diffeomorphisms on the base. For $m\geq 1$, we have $\pi_{0}(\operatorname{Map}(\mathbb{S}^{m},\mathrm{O}(k)))=\pi_{m}(\mathrm{O}(k))\rtimes\mathbb{Z}/2$, so as $\pi_{m}(O(2))=\pi_{m}(\mathbb{S}^{1})=0$ for $m\geq 2$ and $\pi_{1}(O(m))=\mathbb{Z}/2$ for $m\geq 3$, we see that the two mapping spaces in (4.1) have finitely many components. Since $\pi_{0}(\operatorname{Diff}(\mathbb{S}^{m}))$ is finite for $m\neq 4$ by Corollary 5.2 below, it follows from the long exact sequence in homotopy groups that $\pi_{0}(\operatorname{Isom}_{b}(D^{\pm}))$ and thus $H_{\pm}$ is finite for $n\neq 6$. For $n=6$, the group $\pi_{0}(\operatorname{Diff}(\mathbb{S}^{n-2}))$ maybe be infinite, but since $\pi_{0}(\operatorname{Map}(\mathbb{S}^{n-2},O(2)))$ is finite even for $n=6$ and $\partial D^{\pm}=\mathbb{S}^{n-2}\times\mathbb{S}^{1}$, for finiteness of $H_{+}$ it actually suffices that that the image of $(-)\times\mathbb{S}^{1}\colon\pi_{0}(\operatorname{Diff}(\mathbb{S}^{n-2}))\rightarrow\pi_{0}(\operatorname{Diff}(\mathbb{S}^{n-2}\times\mathbb{S}^{1}))$ is finite for $n=6$ which holds by Lemma 5.3 below. ∎
@@
The space of leaves $X_{m}$ of $\mathcal{F}_{m}$ is canonically isometric to the quotient of $\tilde{M}$ by the action of the group $\mathbb{(}\mathbb{Z}\times H_{1})\cdot\mathcal{H}_{m}$. Hence, $X_{m}$ is the circle of length $\frac{1}{m}$. Thus, the leaves of $\mathcal{F}_{m}$ are fibers of a Riemannian submersion $P_{m}:M\to X_{m}$. As explained above, its fibers have fundamental groups $H_{m}$ and constant mean curvature, and this mean curvature does not depend on the fiber. Taking the fiber with minimal volume, we deduce that at this fiber the mean curvature is 0, thus all fibers of $P_{m}$ are minimal. Composing $P_{m}$ with an isoparametric function $g_{m}:X_{m}\to\mathbb{R}$, we obtain an isoparametric function $f_{m}:=g_{m}\circ P_{m}:M\to\mathbb{R}$ by [GT13, Proposition 2.4]. The isoparametric foliation induced by $f_{m}$ is exactly $\mathcal{F}_{m}$. But the fundamental groups of the fibers of $\mathcal{F}_{m}$ are different, for different $m$. This implies the claim for $\dim(M)=4$ and thus also in higher dimensions by taking products with a torus. ∎
@@ 5. Topological background
In this final section, we collect the differential-topological results used in the previous sections. To high-dimensional topologists, essentially all of them are either clear or well-known; we merely explain how to extract them from the literature for more geometrically minded readers. We adopt the notation from Section 2.3. In particular, for closed manifolds $M$ we write $\operatorname{Diff}(M)=\operatorname{Diff}(M,M)$ and $\operatorname{Homeo}(M)=\operatorname{Homeo}(M,M)$ for the topological groups of diffeomorphisms or homeomorphisms of $M$. If $M$ is compact but may have boundary, we require diffeomorphisms or homeomorphisms to agree with the identity in a neighbourhood of the boundary, which we indicate by a $\partial$-subscript. If we restrict to orientation-preserving diffeomorphisms or homeomorphisms, we add a ++-superscript.
@@
By the parametrised isotopy extension sequence (see e.g. [Wal16, Theorem 6.1.1]), restriction along $D^{n}_{-}\subset\mathbb{S}^{n}$ yields a fibration $\operatorname{Diff}^{+}(\mathbb{S}^{n})\rightarrow\mathrm{Emb}^{+}(D^{n}_{-},\mathbb{S}^{n})$ to the space of orientation-preserving embeddings $D^{n}_{-}\hookrightarrow\mathbb{S}^{n}$. The fiber over the inclusion is $\operatorname{Diff}_{\partial}(D_{+}^{n})$. Taking derivatives induces a homotopy equivalence $\mathrm{Emb}^{+}(D^{n}_{-},\mathbb{S}^{n})\simeq\mathrm{GL}^{+}_{n+1}(\mathbb{R})\simeq\mathrm{SO}(n+1)$, so the standard $\mathrm{SO}(n+1)$-action on $\mathbb{S}^{n}$ yields a section of $\operatorname{Diff}^{+}(\mathbb{S}^{n})\rightarrow\mathrm{Emb}^{+}(D^{n}_{-},\mathbb{S}^{n})$ up to homotopy. The long exact sequence in homotopy groups then implies that $\mathrm{ext}_{n}$ is an isomorphism, so $\smash{\widetilde{\mathrm{ext}}_{n}}$ is surjective too. To show injectivity of $\widetilde{\mathrm{ext}}_{n}$, we prove that $\widetilde{\mathrm{ext}}_{n}\circ\mathrm{glue}_{n}$ is injective: let $\varphi]$ be a class in the kernel, then there is an orientation-preserving diffeomorphism
$\hskip-14.22636pt\smash{D^{n+1}_{+}\cup_{\mathrm{glue}(\varphi)}(\mathbb{S}^{n}\times[0,1])\cup_{\mathbb{S}^{n}\times\{1\}}D^{n+1}_{-}\overset{\Phi}{\cong}D^{n+1}_{+}\cup_{\mathrm{id}}(\mathbb{S}^{n}\times[0,1])\cup_{\mathbb{S}^{n}\times\{1\}}D^{n+1}_{-}.}$
@@
Since $\mathrm{ext}_{4}$ from (5.1) is an isomorphism by Lemma 5.1, it suffices to show that the composition $\pi_{0}(\operatorname{Diff}_{\partial}(D_{+}^{4}))\rightarrow\pi_{0}(\operatorname{Diff}_{\partial}(D_{+}^{4}\times\mathbb{S}^{1}))\rightarrow\pi_{0}(\operatorname{Diff}^{+}(\mathbb{S}^{4}\times\mathbb{S}^{1}))$ is trivial. We will see that this is already the case for the first map. Fix a basepoint $\in D^{1}_{+}\subset\mathbb{S}^{1}$ and consider the restriction map $\pi_{0}(\operatorname{Diff}_{\partial}(D_{+}^{4}\times\mathbb{S}^{1}))\rightarrow\pi_{0}(\mathrm{Emb}_{\partial}(D_{+}^{4}\times*,D_{+}^{4}\times\mathbb{S}^{1}))$ to the set of isotopy classes of embeddings that agree with the inclusion $\mathrm{inc}\colon D_{+}^{4}\times*\subset D_{+}^{4}\times\mathbb{S}^{1}$ near the boundary. By isotopy extension, the fibers of this map are either empty or admit a surjection from $\pi_{0}(\operatorname{Diff}_{\partial}(D_{+}^{4}\times D^{1}_{-}))\cong\pi_{0}(\operatorname{Diff}_{\partial}(D^{5}))$ which is trivial by Lemma 5.1, so $\pi_{0}(\operatorname{Diff}_{\partial}(D_{+}^{4}\times\mathbb{S}^{1}))\rightarrow\pi_{0}(\mathrm{Emb}_{\partial}(D_{+}^{4}\times*,D_{+}^{4}\times\mathbb{S}^{1}))$ is injective. It thus suffices to show that the map $\pi_{0}(\operatorname{Diff}_{\partial}(D_{+}^{4}))\rightarrow\pi_{0}(\mathrm{Emb}_{\partial}(D_{+}^{4}\times*,D_{+}^{4}\times\mathbb{S}^{1}))$ which sends $\varphi]\in\pi_{0}(\operatorname{Diff}_{\partial}(D_{+}^{4}))$ to $\mathrm{inc}\circ\varphi]$ is constant. By identifying $D_{+}^{1}\subset\mathbb{S}^{1}$ with $[0,1]$ one sees that this maps sends concordant diffeomorphisms to the same class of embeddings, so it factors over the group $\widetilde{\pi}_{0}(\operatorname{Diff}_{\partial}(D_{+}^{4}))$. The latter is trivial by Lemma 5.1, so the claim follows. ∎